OpenStudy (anonymous):

-6(n-9)= 48 What are thee steps for this problemm ??

OpenStudy (anonymous):

distribute

OpenStudy (anonymous):

distribute -6 to n and 9

OpenStudy (anonymous):

distribute the -6

OpenStudy (ybarrap):

-6 (n-9) = 48 Divide both sides by -6: n-9 = -8 Add 9 to both sides: n = 1

OpenStudy (anonymous):

that equals -6n+54=48

OpenStudy (anonymous):

-6(n-9)=48 Divide each term in the equation by 6. -(6(n-9))/(6)=(48)/(6) Simplify the left-hand side of the equation by canceling the common factors. -n+9=(48)/(6) Simplify the right-hand side of the equation by simplifying each term. -n+9=8 Since 9 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 9 from both sides. -n=-9+8 Add 8 to -9 to get -1. -n=-1 Multiply each term in the equation by -1. -n*-1=-1*-1 Multiply -n by -1 to get n. n=-1*-1 Multiply -1 by -1 to get 1. n=1

OpenStudy (anonymous):

Is it different for addition? likee 5(6+r)= -70

OpenStudy (anonymous):

no same

OpenStudy (anonymous):

5(6+r)=-70 Divide each term in the equation by 5. (5(6+r))/(5)=-(70)/(5) Simplify the left-hand side of the equation by canceling the common factors. r+6=-(70)/(5) Simplify the right-hand side of the equation by simplifying each term. r+6=-14 Since 6 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 6 from both sides. r=-6-14 Subtract 14 from -6 to get -20. r=-20

OpenStudy (anonymous):

soo it would bee likee get rid of the 5?

OpenStudy (anonymous):

ur objective is to find out what r is

OpenStudy (anonymous):

okk.. soo would i didvide 5? or times it? or do it to the 6?

OpenStudy (anonymous):

times the 5 to 6

OpenStudy (anonymous):

then times the 5 to r....to make it disappear